# Statistical Multiport-Network Modeling and Efficient Discrete Optimization of RIS

### v1

@@ Abstract
This Letter fills the research gap on physics-consistent optimization for reconfigurable intelligent surfaces (RISs) with mutual coupling (MC) and 1-bit-tunable elements, a common hardware constraint in existing RIS prototypes. We compare a model-based method (temperature-annealed back-propagation) and model-agnostic methods (coordinate descent, genetic algorithm), and evaluate potential benefits of intelligently initializing these methods. To facilitate our evaluation, we introduce a technique for generating statistical ensembles of multiport-network model parameters, wherein a single hyper-parameter adjusts the MC strength. The technique is a generalization of Rayleigh fading to radio environments with deterministic programmability, and it accounts for passivity constraints as well as the coherent-backscattering effect. We find that, except when MC is negligible, coordinate descent with random initialization yields the most favorable trade-off in terms of performance, execution time and memory usage. We expect our findings to extend to beyond-diagonal RIS, stacked intelligent metasurfaces, dynamic metasurface antennas, and wave-domain physical neural networks.
Index Terms:
Reconfigurable intelligent surface, multi-port network theory, 1-bit programmability, hardware constraint, coherent backscattering, multi-fidelity optimization.
@@ I Introduction
An MC-aware RIS optimization problem is generally high-dimensional, non-convex, and non-linear; hardware constraints like 1-bit tunable RIS elements further complicate the optimization. Closed-form solutions only exist for RIS with tunable inter-element coupling and without any hardware constraints [3]. Recent theoretical studies on MC-aware RIS optimization are model-based (i.e., using a physics-consistent system model whose parameters are assumed to be known) and assume continuously tunable RIS elements. Incremental RIS configuration updates are at the core of these methods, either because utilized Neumann series approximations inherently require sufficiently small configuration updates [4, 5, 6, 7, 8], or because techniques like gradient-descent are used [9, 10, 11]. Therefore, none of these approaches can be straightforwardly applied to an RIS with 1-bit programmable elements. Moreover, many of these theoretical studies simplify physics-consistent models from multi-port network theory (MNT) by assuming unilateral propagation [12], neglecting the MC between antennas and/or the MC between RIS elements originating from multi-bounce paths in the radio environment.
@@
To evaluate optimization techniques in theoretical work, physical-model parameters are required. Existing theoretical works on physics-consistent RIS optimization are fundamentally limited by the difficulty of generating deterministic channel realizations. On the one hand, theoretical studies require analytical tractability such that they can only explore very simple free-space scenarios without structural scattering [4], sometimes adding fading effects ad hoc without considering their contribution to MC. On the other hand, end-to-end channel evaluations are costly in full-wave numerical studies of more complex scenarios; estimating model parameters of more complex scenarios is also quite costly in full-wave simulations [15, 13, 16] or experiments [17, 18, 14, 19, 20, 21]. A method to generate statistical ensembles of physical-model parameters, ideally with controllable MC strength, is missing.
Our contributions are summarized as follows. First, we introduce a simple method to generate statistical realizations of MNT parameters for rich-scattering RIS-parametrized radio environments with a single hyper-parameter to adjust the MC strength between RIS elements. Second, using this technique, we systematically compare model-agnostic and model-based strategies to optimize 1-bit programmable RISs with different levels of MC strength in rich-scattering environments; we evaluate performance, execution time and memory usage. Where appropriate, we explore benefits of intelligently initializing the methods based on a dictionary search or a simplified affine model.
Notation: $\mathbf{A}_{\mathcal{BC}}$ denotes the block of the matrix $\mathbf{A}$ whose row [column] indices are in the set $\mathcal{B}$ [$\mathcal{C}$]. $\mathrm{diag}(\mathbf{a})$ denotes the matrix whose diagonal entries are those of the vector $\mathbf{a}$ and all other entries are zero. $\mathbf{I}_{a}$ denotes the $a\times a$ identity matrix.
@@ II System Model
In this section, we first describe the mathematical structure of the well-established MNT system model for RIS-parametrized radio environments and relate it to simplified models; then, we explain our novel technique for generating statistical ensembles of the MNT model parameters with controllable MC strength.
@@
II-A High-Fidelity Physics-Compliant MNT Model
To apply MNT as shown in Fig. 1, we assume that the antenna ports are lumped and that the RIS elements’ programmability originates from tunable lumped elements (e.g., PIN diodes). A tunable lumped element can be described by an auxiliary lumped port terminated by a tunable load. With $N_{\mathrm{T}}$ transmitting antennas, $N_{\mathrm{R}}$ receiving antennas, and $N_{\mathrm{S}}$ RIS elements, the static parts of the RIS-parametrized radio environment are compactly characterized as a time-invariant, linear, passive, reciprocal $N$-port network, where $N=N_{\mathrm{T}}+N_{\mathrm{R}}+N_{\mathrm{S}}$. The corresponding scattering matrix $\tilde{\mathbf{S}}\in\mathbb{C}^{N\times N}$ incorporates the effects of all scattering objects in the radio environment (including structural scattering of the antennas and RIS elements) [22, 17, 13]. For simplicity, we use the same reference impedance $Z_{0}$ at all ports to define $\tilde{\mathbf{S}}$, and we assume that the signal generators and detectors attached to the transmitting and receiving antennas, respectively, are matched to $Z_{0}$. Throughout this Letter, we do not make any further simplifying assumptions such as the unilateral approximation or neglecting certain coupling effects.
Within MNT, the end-to-end wireless channel matrix $\mathbf{H}\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ is defined as [7, 14, 23]
@@ II-A High-Fidelity Physics-Compliant MNT Model
where $\mathcal{R}$, $\mathcal{T}$, and $\mathcal{S}$ denote the sets of port indices associated with receiving antennas, transmitting antennas, and RIS elements, respectively; $\mathbf{\Phi}=\mathrm{diag}(\mathbf{c})\in\mathbb{C}^{N_{\mathrm{S}}\times N_{\mathrm{S}}}$, where $\mathbf{c}=[r_{1},r_{2},\dots,r_{N_{\mathrm{S}}}]\in\mathbb{C}^{N_{\mathrm{S}}}$ characterizes the reconfigurable parts of the RIS-parametrized radio environment (i.e., the RIS elements’ tunable lumped components), $r_{i}$ being the reflection coefficient of the tunable load terminating the auxiliary port associated with the $i$th RIS element.
@@
[…]
II-B Low-Fidelity Affine Models
@@
where passivity guarantees convergence [14]. The $k$th term of the series captures all paths involving $k$ bounces between RIS elements [2, 14]. The spectral radius $R$ (i.e., the largest eigenvalue magnitude) of $\mathbf{\Phi}\tilde{\mathbf{S}}_{\mathcal{{SS}}}$ governs the convergence speed of the series in (3). The series collapses to a single term only if all entries of $\tilde{\mathbf{S}}_{\mathcal{{SS}}}$ vanish, in which case one recovers the widespread simplified CASC model with an affine mapping
@@
from $\mathbf{\Phi}$ to $\mathbf{H}$. Importantly, zero MC alone does not guarantee this collapse of the series. Indeed, if the diagonal entries of $\tilde{\mathbf{S}}_{\mathcal{{SS}}}$ do not vanish, the series does not collapse. Physically, the series represents bounces confined to the interfaces between radio environment and load network in such cases, i.e., not between different auxiliary lumped ports. These bounces at the $i$th auxiliary port arise because the radio environment does not present an impedance of $Z_{0}$ to the $i$th load. Hence, the value of $R$ is generally not directly indicative of the MC strength.
@@
II-C Statistical MNT Model Parameter Generation
Having described and analyzed the mathematical structure of the MNT system model, we now discuss the choice of MNT-model parameters which are the allowed values of $r_{i}$ and the entries of $\tilde{\mathbf{S}}$. Regarding the former, most theoretical studies assume that $r_{i}$ can take any complex value whose magnitude does not exceed unity (due to passivity), or that $r_{i}$ can be continuously tuned based on a model that intertwines phase and amplitude response [25]. In contrast, given our focus on 1-bit programmability, we work with $r_{i}\in\{-1,1\}$ in this Letter. Regarding the latter, we now introduce a novel technique for generating statistical ensembles of $\tilde{\mathbf{S}}$ that emulate rich-scattering environments with tunable MC between RIS elements. This is in contrast to deterministic approaches that evaluate $\tilde{\mathbf{S}}$ for a concrete setup, e.g., using a discrete-dipole approximation [26, 27], for which the generation of large ensembles of $\tilde{\mathbf{S}}$ is prohibitively costly.
Our technique is inspired by the conventional generation of Rayleigh fading channel realizations for scenarios without RIS. Therein, the channel coefficients are drawn from a circularly symmetric zero-mean complex Gaussian distribution. These ensemble statistics match those expected in an ideal mode-stirred reverberation chamber with at least half-wavelength antenna spacing [28, 29, 30]. Moreover, it is known that in this setting the variance of the reflection coefficients is twice that of the transmission coefficients due to coherent backscattering [31]. As evident in the MNT model, RIS elements are nothing but antennas terminated by tunable loads. Thus, $\tilde{\mathbf{S}}$ corresponds to the scattering matrix of an ideal reverberation chamber containing $N$ antennas with at least half-wavelength spacing. The half-wavelength spacing between RIS elements is common in most RIS prototypes [1]. We hence draw the diagonal and off-diagonal entries of $\tilde{\mathbf{S}}$ from circularly symmetric zero-mean complex Gaussian distributions differing by a factor of two in variance. We impose that $\tilde{\mathbf{S}}$ is symmetric to respect reciprocity, and we scale all entries of $\tilde{\mathbf{S}}$ by a factor of 1/15 to respect passivity (i.e., to ensure that all singular values of $\tilde{\mathbf{S}}$ are well below unity). Finally, to adjust the MC strength between the RIS elements, we introduce a positive, real hyper-parameter $\kappa$ that scales all off-diagonal entries of $\tilde{\mathbf{S}}_{\mathcal{SS}}$. The value of $\mu_{\mathrm{n}}$ scales by definition with the same factor $\kappa$. We impose an upper limit on $\kappa$ to respect the passivity of $\tilde{\mathbf{S}}$.
@@ III-2 Coordinate Descent (CD)
1 Initialize RIS configuration: $\mathbf{c}_{\mathrm{curr}}\leftarrow\mathbf{c}_{\mathrm{init}}$.
@@ 5  $j\leftarrow(j\ \mathrm{mod}\ N_{\mathrm{S}})+1$
Output: Optimized RIS configuration $c_{\mathrm{curr}}.$
@@ III-3 Temperature-Annealed Back-Propagation (TABP)
TABP optimizes a relaxed continuous version of the binary problem using gradient descent, gradually driving the distribution toward a binary one [32, 33]. Specifically, a continuous, real-valued variable $\mathbf{z}\in\mathbb{R}^{N_{\mathrm{S}}}$ is mapped to $\tilde{\mathbf{c}}$ using the temperature-scaled $\tanh$ function: $\tilde{\mathbf{c}}(\mathbf{z},t)=c^{\prime}+\frac{1}{2}\bigl{(}1+\tanh(\mathbf{z}/t)\bigr{)}(c^{\prime\prime}-c^{\prime})$, where $c^{\prime}$ and $c^{\prime\prime}$ are the two allowed values for the entries of $\mathbf{c}$, and $t$ is the temperature parameter. As the epochs advance, $t$ decreases and the distribution of the entries of $\tilde{\mathbf{c}}$ becomes more and more binary. After training, we enforce the binarization: $\mathbf{c}_{\mathrm{curr}}=\lim_{t\to 0}\tilde{\mathbf{c}}(\mathbf{z},t)$. We implement the TABP Algorithm 2 based on the CASC and MNT models, using the Adam optimizer [34] with default hyper-parameters to minimize $\mathcal{C}(\mathbf{H})$; we consider the same initializations as for CASC-CD and MNT-CD. We train for up to $e_{\mathrm{max}}=400$ epochs with $\epsilon=10^{-4}$, $t_{\mathrm{start}}=1$, and $t_{\mathrm{end}}=0.1$.
1 $\mathbf{z}\;\leftarrow\;t_{\rm start}\,\operatorname{atanh}((1-\epsilon)(2(\mathbf{c}_{\rm init}-c^{\prime})/(c^{\prime\prime}-c^{\prime})-1))$.
2 $\tilde{\mathbf{c}}\leftarrow c^{\prime}+\tfrac{1}{2}\bigl{(}1+\tanh(\mathbf{z}/t_{\rm start})\bigr{)}\,(c^{\prime\prime}-c^{\prime})$.
3 $\mathcal{C}_{\mathrm{curr}}\leftarrow\mathcal{C}(\mathbf{H}(\tilde{\mathbf{c}}))$.
@@
5  $t\leftarrow t_{\text{start}}\left(t_{\text{end}}/t_{\text{start}}\right)^{e/e_{\mathrm{max}}}$.
6  $\tilde{\mathbf{c}}\leftarrow c^{\prime}+\frac{1}{2}\bigl{(}1+\tanh(\mathbf{z}/t)\bigr{)}(c^{\prime\prime}-c^{\prime})$.
7  $\mathcal{C}_{\mathrm{new}}\leftarrow\mathcal{C}(\mathbf{H}(\tilde{\mathbf{c}}))$.
@@
Output: Optimized RIS configuration $c_{\mathrm{curr}}.$

### v2

@@ Abstract
This Letter addresses the physics-consistent optimization of reconfigurable intelligent surfaces (RISs) with mutual coupling (MC) and 1-bit-programmable RIS elements. This combination of constraints is typical of current prototypes but unexplored in theoretical work. First, we present a simple statistical generator for multiport-network-theory (MNT) parameters of rich-scattering, RIS-parametrized channels. We account for reciprocity, passivity, and coherent backscattering; then, we add a simple hyper-parameter to control the MC strength. Second, we benchmark model-agnostic (dictionary search, coordinate descent, genetic algorithm) and model-based (temperature-annealed back-propagation) strategies under varying MC, with and without intelligent initialization. Except when MC is negligible, coordinate descent with random initialization offers the best trade-off in performance, runtime, and memory. Our insights can guide wireless practitioners who optimize RIS prototypes and other reconfigurable wave systems.
@@ I Introduction
An MC-aware RIS optimization problem is generally high-dimensional, non-convex, and non-linear; hardware constraints like 1-bit tunable RIS elements further complicate the optimization. Closed-form solutions only exist for RIS with tunable inter-element coupling and without any hardware constraints [3]. Recent theoretical studies on MC-aware RIS optimization are model-based (i.e., using a physics-consistent system model whose parameters are assumed to be known) and assume continuously tunable RIS elements. Incremental RIS configuration updates are at the core of these methods, either because utilized Neumann-series approximations inherently require sufficiently small configuration updates [4, 5, 6, 7, 8], or because techniques like gradient-descent are used [9, 10, 11]. Therefore, none of these approaches can be straightforwardly applied to an RIS with 1-bit-programmable elements. Moreover, many of these theoretical studies simplify physics-consistent models from multi-port network theory (MNT) by assuming unilateral propagation [12], neglecting the MC between antennas and/or the MC between RIS elements originating from multi-bounce paths in the radio environment.
@@
Wireless practitioners urgently need guidance on which algorithm to choose to work with current RIS prototypes that are subject to MC and 1-bit-programmability; yet, to our knowledge, no systematic study comparing possible algorithms exists for discrete, MC-aware RIS optimization.
Moreover, existing theoretical works on physics-consistent RIS optimization are fundamentally limited by the difficulty of generating deterministic channel realizations. On the one hand, theoretical studies require analytical tractability such that they can only explore very simple free-space scenarios without structural scattering [4], sometimes adding fading effects ad hoc without considering their contribution to MC. On the other hand, end-to-end channel evaluations are costly in full-wave numerical studies of more complex scenarios; estimating model parameters of more complex scenarios is also quite costly in full-wave simulations [13, 15] or experiments [16, 17, 14, 18, 19, 20]. A method to generate statistical ensembles of physical-model parameters, ideally with controllable MC strength, is missing. A statistical analysis for rich-scattering environments is desirable because MC generally depends on the specific environment [2]; ensemble-averaged analysis provides robust, generalizable insights beyond any single realization.
Our contributions are summarized as follows. First, we propose a simple reverberation-chamber-inspired statistical generator for MNT parameters of rich-scattering, RIS-parametrized radio environments, with a single hyper-parameter to adjust the MC strength between RIS elements. Second, using this technique, we systematically compare model-agnostic and model-based strategies to optimize 1-bit-programmable RISs with different levels of MC strength in rich-scattering environments; we evaluate performance, execution time and memory usage. Where appropriate, we explore benefits of intelligently initializing the methods based on a dictionary search or a simplified affine model.
Notation: $\mathbf{A}_{\mathcal{BC}}$ denotes the block of the matrix $\mathbf{A}$ whose row [column] indices are in the set $\mathcal{B}$ $\mathcal{C}$]. $\mathrm{diag}(\mathbf{a})$ denotes the matrix whose diagonal entries are those of the vector $\mathbf{a}$ and all other entries are zero. $\mathbf{I}_{a}$ denotes the $a\times a$ identity matrix.
@@ II System Model
In this section, we first describe the mathematical structure of the well-established MNT system model for RIS-parametrized radio environments and relate it to simplified models; then, we explain our simple technique for generating statistical ensembles of the MNT model parameters with controllable MC strength.
@@
II-1 High-Fidelity Physics-Consistent MNT Model
To apply MNT as shown in Fig. 1, we assume that the antenna ports are lumped and that the RIS elements’ programmability originates from tunable lumped elements (e.g., PIN diodes). A tunable lumped element can be described by an auxiliary lumped port terminated by a tunable load. With $N_{\mathrm{T}}$ transmitting antennas, $N_{\mathrm{R}}$ receiving antennas, and $N_{\mathrm{S}}$ RIS elements, the static parts of the RIS-parametrized radio environment are compactly characterized as a time-invariant, linear, passive, reciprocal $N$-port network, where $N=N_{\mathrm{T}}+N_{\mathrm{R}}+N_{\mathrm{S}}$. The corresponding scattering matrix $\tilde{\mathbf{S}}\in\mathbb{C}^{N\times N}$ incorporates the effects of all scattering objects in the radio environment (including structural scattering of the antennas and RIS elements) [21, 16, 13]. For simplicity, we use the same reference impedance $Z_{0}$ at all ports to define $\tilde{\mathbf{S}}$, and we assume that the signal generators and detectors attached to the transmitting and receiving antennas, respectively, are matched to $Z_{0}$.
The end-to-end wireless channel matrix $\mathbf{H}\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ can be derived from MNT [7, 8, 22], yielding
@@ II-1 High-Fidelity Physics-Consistent MNT Model
where $\mathcal{R}$, $\mathcal{T}$, and $\mathcal{S}$ denote the sets of port indices associated with receiving antennas, transmitting antennas, and RIS elements, respectively; $\mathbf{\Phi}=\mathrm{diag}(\mathbf{c})\in\mathbb{C}^{N_{\mathrm{S}}\times N_{\mathrm{S}}}$, where $\mathbf{c}=[c_{1},c_{2},\dots,c_{N_{\mathrm{S}}}]\in\mathbb{C}^{N_{\mathrm{S}}}$ characterizes the reconfigurable parts of the RIS-parametrized radio environment (i.e., the RIS elements’ tunable lumped components), $c_{i}$ being the reflection coefficient of the tunable load terminating the auxiliary port associated with the $i$th RIS element.
@@
where $\langle\cdot\rangle$ denotes the ensemble averaging over random RIS configurations so that we have a configuration-agnostic metric. [^1] We average over 100 random RIS configurations to ensure a small standard error of the mean while keeping runtime reasonable.
II-2 Low-Fidelity Affine Models
@@
where passivity guarantees convergence [14]. The $k$th term of the series captures all paths involving $k$ bounces between RIS elements [2, 14]. The spectral radius $R$ (i.e., the largest eigenvalue magnitude) of $\mathbf{\Phi}\tilde{\mathbf{S}}_{\mathcal{{SS}}}$ governs the convergence speed of the series in (3). The series collapses to a single term only if all entries of $\tilde{\mathbf{S}}_{\mathcal{{SS}}}$ vanish, in which case one recovers the widespread simplified cascaded (CASC) model with an affine mapping
@@
from $\mathbf{\Phi}$ to $\mathbf{H}$. Importantly, zero MC alone does not guarantee this collapse of the series. Indeed, if the diagonal entries of $\tilde{\mathbf{S}}_{\mathcal{{SS}}}$ (i.e., the reflection coefficients seen at the auxiliary ports with all other ports matched to $Z_{0}$) do not vanish, the series does not collapse. Physically, the series represents bounces confined to the interfaces between radio environment and load network in such cases, i.e., not between different auxiliary lumped ports. These bounces at the $i$th auxiliary port arise because the radio environment does not present an impedance of $Z_{0}$ to the $i$th load. Hence, the value of $R$ is generally not directly indicative of the MC strength.
@@
II-3 Statistical MNT Model Parameter Generation
Having described and analyzed the mathematical structure of the MNT system model, we now discuss the choice of MNT-model parameters which are the allowed values of $c_{i}$ and the entries of $\tilde{\mathbf{S}}$. Regarding the former, most theoretical studies assume that $c_{i}$ can take any complex value whose magnitude does not exceed unity (due to passivity), or that $c_{i}$ can be continuously tuned based on a model that intertwines phase and amplitude response [24]. In contrast, given our focus on 1-bit programmability, we work with $c_{i}\in\{-1,1\}$ in this Letter. Regarding the latter, we now introduce a simple technique for generating statistical ensembles of $\tilde{\mathbf{S}}$ that emulate rich-scattering environments with tunable MC between RIS elements. This is in contrast to deterministic approaches that evaluate $\tilde{\mathbf{S}}$ for a concrete setup, e.g., using a discrete-dipole approximation [25, 26], for which the generation of large ensembles of $\tilde{\mathbf{S}}$ is prohibitively costly.
Our technique builds upon the statistical description of scattering in reverberation chambers (RCs) [27, 28, 29]. In an ideal mode-stirred RC with at least half-wavelength antenna spacing, the channel coefficients between antenna ports (which are off-diagonal entries of a scattering matrix) are drawn from a circularly symmetric zero-mean complex Gaussian distribution (Rayleigh fading). Moreover, in this setting the variance of the reflection coefficients is twice that of the transmission coefficients due to coherent backscattering [30]. To map these insights to our system model, we recall that RIS elements are antennas terminated by tunable loads. Thus, $\tilde{\mathbf{S}}$ corresponds to the scattering matrix of an ideal reverberation chamber containing $N$ antennas with at least half-wavelength spacing. The half-wavelength spacing between adjacent RIS elements is common in most RIS prototypes [1]. We hence draw the diagonal and off-diagonal entries of $\tilde{\mathbf{S}}$ from circularly symmetric zero-mean complex Gaussian distributions with variance 1 and 0.5, respectively. We impose that $\tilde{\mathbf{S}}$ is symmetric to respect reciprocity. To respect passivity, we scale all entries of $\tilde{\mathbf{S}}$ by a factor $\alpha$ and subsequently discard any realizations of $\tilde{\mathbf{S}}$ that have singular values above unity; we heuristically chose $\alpha=1/15$ but we checked that other choices of $\alpha$ that ensure passivity lead to the same qualitative behavior of our results.
Finally, we introduce ad hoc a simple single-parameter tuning knob to adjust the RIS elements’ MC strength: a positive, real hyper-parameter $\kappa$ scales all off-diagonal entries of $\tilde{\mathbf{S}}_{\mathcal{SS}}$. The value of $\mu_{\mathrm{n}}$ scales by definition with the same factor $\kappa$; we impose an upper limit on $\kappa$ to respect the passivity of $\tilde{\mathbf{S}}$. The simplicity of this ad hoc MC strength control, not requiring the incorporation of detailed geometrical information, justifies its use to study algorithmic performance without committing to any particular hardware layout.
@@ III-2 Coordinate Descent (CD)
1Initialize RIS configuration: $\mathbf{c}_{\mathrm{curr}}\leftarrow\mathbf{c}_{\mathrm{init}}$.
@@ 5  $j\leftarrow(j\ \mathrm{mod}\ N_{\mathrm{S}})+1$
Output: Optimized RIS configuration $\mathbf{c}_{\mathrm{curr}}.$
@@ III-3 Temperature-Annealed Back-Propagation (TABP)
TABP optimizes a relaxed continuous version of the binary problem using gradient descent, gradually driving the distribution toward a binary one [31, 32]. Specifically, a continuous, real-valued variable $\mathbf{z}\in\mathbb{R}^{N_{\mathrm{S}}}$ is mapped to $\tilde{\mathbf{c}}$ using the temperature-scaled $\tanh$ function: $\tilde{\mathbf{c}}(\mathbf{z},t)=c^{\prime}+\frac{1}{2}\bigl(1+\tanh(\frac{\mathbf{z}}{t})\bigr)(c^{\prime\prime}-c^{\prime})$, where $c^{\prime}$ and $c^{\prime\prime}$ are the two allowed values for the entries of $\mathbf{c}$, and $t$ is the temperature parameter. As the epochs advance, $t$ decreases and the distribution of the entries of $\tilde{\mathbf{c}}$ becomes more and more binary. After training, we enforce the binarization: $\mathbf{c}_{\mathrm{curr}}=\lim_{t\to 0}\tilde{\mathbf{c}}(\mathbf{z},t)$. We implement the TABP Algorithm 2 based on the CASC and MNT models, using the Adam optimizer [33] with default hyper-parameters to minimize $\mathcal{C}(\mathbf{H})$; we consider the same initializations as for CASC-CD and MNT-CD. We train for up to $e_{\mathrm{max}}=200$ epochs with $\epsilon=10^{-6}$, $t_{\mathrm{start}}=1$, and $t_{\mathrm{end}}=0.1$.
[…]
2 $\tilde{\mathbf{c}}\leftarrow c^{\prime}+\tfrac{1}{2}\bigl(1+\tanh\!\bigl(\tfrac{\mathbf{z}}{t_{\rm start}}\bigr)\bigr)(c^{\prime\prime}-c^{\prime})$.
3 $)\mathcal{C}_{\mathrm{curr}}\leftarrow\mathcal{C}(\mathbf{H}(\tilde{\mathbf{c}}$)).
@@
5  $t\leftarrow t_{\text{start}}\left(\tfrac{t_{\text{end}}}{t_{\text{start}}}\right)^{e/e_{\mathrm{max}}}$.
6  $\tilde{\mathbf{c}}\leftarrow c^{\prime}+\frac{1}{2}\bigl(1+\tanh(\frac{\mathbf{z}}{t})\bigr)(c^{\prime\prime}-c^{\prime})$.
7  $)\mathcal{C}_{\mathrm{new}}\leftarrow\mathcal{C}(\mathbf{H}(\tilde{\mathbf{c}}$)).
@@
Output: Optimized RIS configuration $\mathbf{c}_{\mathrm{curr}}.$